very difficult the ans found here wrong

A sprinter runs a 100 m dash in 12.0 s. She starts from rest

with a constant acceleration ax for 3.0 s and then runs with constant

speed for the remainder of the race. What is the value of ax?
Answer #1

ax=3.2m/s2

answered by: Kelsie Johnson
Answer #2

SOLUTION :


!00m completed in 12 sec.


Initial speed = 0 m/s

Run for 3 sec at constant acceleration ax and then runs at constant speed.


Speed after 3 sec = ax * 3  = 3 ax m/s


Average speed during first 3 sec of acceleration = (3 ax + 0)/2 = 1.5 ax m/s 


So, we can consider run of 3 sec. at average speed of 1.5  ax m/s and remaining 9 sec at the speed of 9 sec.


So, 


Distance run in 12 sec = 1.5 ax * 3 + 3.0 ax * 9 

=> 100 = 4.5 ax + 27 ax

=> 100 = 31.5 ax 

=> ax = 100/31.5

=> ax = 3.1746 m/s^2 = 3.175 m/s^2 approx. (ANSWER).

answered by: Tulsiram Garg

> In the 6th line, correct, " remaining 9 sec at the speed of 9 sec. " as " remaining 9 sec at the speed of 3 ax m/s " .

Tulsiram Garg Tue, Dec 14, 2021 8:03 AM